# Semiformal method I. Pólya’s theorem for the complete graph $K_{\mathbb{N}}$ with complex edge weights

### v6

@@ 4 Extensions of Pólya’s theorem with $v\neq 1$
2. If $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}$ is a $v$-transitive 0-convex light weight with $v\in V(h)$, $B_{h},C_{h,v}\in H_{1}$, $(B_{h,v})_{1/\xi}(1)_{1}$ exists and if $(B_{h})_{1/\xi}(1)_{1}=\pm\infty$, then
$\lim_{x\nearrow 1}F_{(A_{h,v})_{1/\xi}}(x)\in\{-1,\,1\}\,.$
Proof. 1. By Proposition 2.31, $D_{h}(x)=\frac{1}{1-\xi x}$. By the subseries bound principle we have
$0\leq\xi^{-n}b_{n}^{h},\,\xi^{-n}b_{n}^{h,v},\,\xi^{-n}c_{n}^{h,v},\,\xi^{-n}e_{n}^{h,v}\leq\xi^{-n}d_{n}^{h}=1$
@@
for every $n\in\mathbb{N}_{0}$. Hence $B_{h}(x)$, $B_{h,v}(x)$, $C_{h,v}(x)$ and $E_{h,v}(x)$ are in $H_{1}$. Any shortest walk joining $1$ and $v$ in $G(h)$ shows that for some $m\in\mathbb{N}$ we have $\xi^{-m}e_{m}^{h,v}>0$. Thus the function $F_{(E_{h,v})_{1/\xi}}(x)$ is positive and increasing on $(0,1)$. Since $F_{(C_{h,v})_{1/\xi}}(x)$ is bounded on $[0,1)$ by item 2 of Proposition 4.1, $F_{(B_{h,v})_{1/\xi}}(x)$ is bounded by item 3. Since by item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-\xi x}$, Proposition 2.15 gives
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}=(C_{h,v})_{1/\xi}(1)_{1}=(C_{h,v})_{1/\xi}(1)_{2}\,.$
By item 2 of Proposition 4.1, Propositions 2.2, 2.4 and 2.5, and by Theorem 2.18 we have
$(C_{h,v}^{2})_{1/\xi}(1)_{1}=\lim_{x\nearrow 1}F_{(C_{h,v}^{2})_{1/\xi}}(x)=1-\frac{\lim_{x\nearrow 1}F_{(B_{h,v})_{1/\xi}}(x)}{\lim_{x\nearrow 1}F_{(B_{h})_{1/\xi}}(x)}=1\,$,}
because $F_{(B_{h,v})_{1/\xi}}(x)$ is bounded on $[0,1)$ and $\lim_{x\nearrow 1}F_{(B_{h})_{1/\xi}}(x)=B_{h}(1)_{1}=+\infty$. By Proposition 2.25, $(C_{h,v})_{1/\xi}(1)_{2}$ exists and
$\big((C_{h,v})_{1/\xi}(1)_{2}\big)^{2}=\big((C_{h,v})_{1/\xi}\big)^{2}(1)_{2}=(C_{h,v}^{2})_{1/\xi}(1)_{2}=(C_{h,v}^{2})_{1/\xi}(1)_{1}=1\,.$
Since $(C_{h,v})_{1/\xi}(1)_{2}\geq 0$, we have $(C_{h,v})_{1/\xi}(1)_{2}=1$ and the stated formula follows.
2. By item 1 of Proposition 4.1 and by Proposition 2.31, $A_{h,v}(x)=C_{h,v}(x)$. By item 2 of Proposition 4.1 and by Theorems 2.16 and 2.17,
$\lim_{x\nearrow 1}F_{(C_{h,v}^{2})_{1/\xi}}(x)=1-\frac{\lim_{x\nearrow 1}F_{(B_{h,v})_{1/\xi}}(x)}{\lim_{x\nearrow 1}F_{(B_{h})_{1/\xi}}(x)}=1-\frac{(B_{h,v})_{1/\xi}(1)_{1}}{\pm\infty}=1\,.$}
By the continuity of the functions $F_{\dots}$,
$\lim_{x\nearrow 1}F_{(A_{h,v})_{1/\xi}}(x)=\lim_{x\nearrow 1}F_{(C_{h,v})_{1/\xi}}(x)\in\{-1,\,1\}\,.$
@@
5 Concluding remarks
In [17] we consider the same problem of determining formulas for quantities such as $A_{h,v}(1)=\sum_{n\geq 0}h(W(v,n))$ in the case of formal weights $h$. These are maps
$h\colon\mathbb{N}_{2}\to\mathbb{C}[[x_{1},\,x_{2},\,\dots,\,x_{k}]]\,$,
where the last integral domain of formal power series in $k$ variables is endowed with the usual non-Archimedean norm.
The main novelty of our generalization is this. In equalities like
$P_{d}(\overline{v},\,n)=h(W(v,\,n))$
that was stated in the Introduction, the left-hand side involves, in the standard rigorous probabilistic treatment, an implicit probability space supporting the corresponding Markov chain. Construction of this probability space is far from trivial; see, for example, [6]. In our generalization, the right-hand side is completely self-contained.

### v8

@@ 4.3 Pólya’s theorems for $v\neq 1$ and $\xi$-convex weights
2. Let $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}$ be a $v$-transitive 0-convex light weight such that $v\in V(h)$. If $B_{h}(x),C_{v,h}(x)\in H_{1}$, $(B_{v,h})_{1/\xi}(1)_{1}$ exists and if $(B_{h})_{1/\xi}(1)_{1}=\pm\infty$, then
$\lim_{x\nearrow 1}(A_{v,\,h})_{1/\xi}(x)\in\{-1,\,1\}\,.$
Proof. 1. By Proposition 2.31, $D_{h}(x)=\frac{1}{1-\xi x}$. By the subseries bound,
$0\leq\xi^{-n}h(B_{n}),\,\xi^{-n}h(B_{v,\,n}),\,\xi^{-n}h(C_{v,\,n}),\,\xi^{-n}h(E_{v,\,n})\leq\xi^{-n}h(D_{n})=1$
@@
for every $n\in\mathbb{N}_{0}$. Hence $B_{h}(x)$, $B_{v,h}(x)$, $C_{v,h}(x)$, and $E_{v,h}(x)$ are in $H_{1}$. Any shortest walk joining $1$ and $v$ in $G(h)$ shows that for some $m\in\mathbb{N}$ we have $\xi^{-m}h(E_{v,m})>0$. Thus the function $(E_{v,h})_{1/\xi}(x)$ is positive and increasing on $(0,1)$. Since $(C_{v,h})_{1/\xi}(x)$ is bounded on $[0,1)$ by item 2 of Proposition 4.2, $(B_{v,h})_{1/\xi}(x)$ is bounded by item 3. Since by item 1 of Proposition 4.2 and by Proposition 2.31 we have $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)=C_{v,h}(x)\frac{1}{1-\xi x}$, Proposition 2.15 gives
$\lim_{n\to\infty}\xi^{-n}h(A_{v,\,n})=(C_{v,\,h})_{1/\xi}(1)_{1}=(C_{v,\,h})_{1/\xi}(1)_{2}\,.$
By item 2 of Proposition 4.2, by the arithmetic of functional limits, and by Theorem 2.18, we have
$((C_{v,\,h})_{1/\xi})^{2}(1)_{1}=\lim_{x\nearrow 1}(C_{v,\,h})_{1/\xi}(x)^{2}=1-\frac{\lim_{x\nearrow 1}(B_{v,\,h})_{1/\xi}(x)}{\lim_{x\nearrow 1}(B_{h})_{1/\xi}(x)}=1\,$,}
because the function $(B_{v,h})_{1/\xi}(x)$ is bounded on $[0,1)$ and $\lim_{x\nearrow 1}(B_{h})_{1/\xi}(x)=B_{h}(1)_{1}=+\infty$. Hence, by nonnegativity, $(C_{v,h})_{1/\xi}(1)_{1}=(C_{v,h})_{1/\xi}(1)_{2}=1$ because
$\big((C_{v,\,h})_{1/\xi}(1)_{2}\big)^{2}=\big((C_{v,\,h})_{1/\xi}\big)^{2}(1)_{2}$
by (the version for nonnegative coefficients of) Proposition 2.20. The stated formula follows.
2. By item 1 of Proposition 4.2 and by Proposition 2.31, $A_{v,h}(x)=C_{v,h}(x)$. By item 2 of Proposition 4.2 and by the arithmetic of functional limits,
$\lim_{x\nearrow 1}(C_{v,\,h})_{1/\xi}(x)^{2}=1-\frac{\lim_{x\nearrow 1}(B_{v,\,h})_{1/\xi}(x)}{\lim_{x\nearrow 1}(B_{h})_{1/\xi}(x)}=1-\frac{(B_{v,\,h})_{1/\xi}(1)_{1}}{\pm\infty}=1-0=1\,.$}
In the second equality we used Theorems 2.16 (in the numerator) and 2.17 (in the denominator). Since the involved real functions are continuous on $[0,1)$, the limit $\lim_{x\nearrow 1}(A_{v,h})_{1/\xi}(x)$ exists and
$\lim_{x\nearrow 1}(A_{v,\,h})_{1/\xi}(x)=\lim_{x\nearrow 1}(C_{v,\,h})_{1/\xi}(x)\in\{-1,\,1\}\,.$
@@
We feel that part 2 of the theorem is not completely satisfactory. Perhaps the assumption that $(B_{v,h})_{1/\xi}(1)_{1}$ exists may be dropped, and it is true that if $(B_{h})_{1/\xi}(1)_{1}=\pm\infty$, then always
$\lim_{x\nearrow 1}\frac{(B_{v,\,h})_{1/\xi}(x)}{(B_{h})_{1/\xi}(x)}=0\,.$}
5 Overview of general Pólya’s theorems
Thm type assumption conclusion
3.2 $v=1,{}_{1}$ $A(1),B(1),C(1),D(1)$ $A(1)=(1-\frac{1}{B(1)})D(1)$
3.3 $v=1,{}_{2}$ $\sum|h(B_{n})|<1,D(1)$ $A(1)=(1-\frac{1}{B(1)})D(1)$
3.4 $v=1,\infty$ $\xi\mathbb{R}_{\geq 0},D(1)=+\infty$ $A(1)=+\infty$
3.5 $v=1,{}_{1}$, $\xi$-c. $\xi\mathbb{R}_{\geq 0}$
$\lim h(A_{n})=1-\frac{1}{B(1)}\in$
$\in(0,1]$
3.6.1 $v=1,{}_{1}$, $\xi$-c. $B(1),C(1)$ $\lim h(A_{n})=1-\frac{1}{B(1)}$
3.6.2 $v=1,{}_{1}$, 0-c. $B(1),C(1)$ $A(1)=1-\frac{1}{B(1)}$
3.7.1 $v=1,{}_{2}$, $\xi$-c. $\sum|h(B_{n})|<1$ $\lim h(A_{n})=1-\frac{1}{B(1)}$
3.7.2 $v=1,{}_{2}$, 0-c. $\sum|h(B_{n})|<1$ $A(1)=1-\frac{1}{B(1)}$
3.8.1 $v=1,\infty$, $\xi$-c. $\xi\mathbb{R}_{\geq 0},B(1)=+\infty$ $\lim h(A_{n})=1$
3.8.2 $v=1,\infty$, 0-c.
$\xi\mathbb{R},B(x),C(x)\in H_{1}$,
$B(1)=+\infty$
$\lim A(x)=1$
4.5 $v\neq 1,{}_{1}$
$A(1),B(1)\neq 0,B_{v}(1)$,
$C_{v}(1),D(1)$
$A(1)\in(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}D(1)$
4.6 $v\neq 1,{}_{1}$
$A(1),B(1),B_{v}(1)$,
$C_{v}(1),D(1),E(1)$
$A(1)\in(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}D(1)$
4.7 $v\neq 1,{}_{2}$
$B_{v}(1),D(1)$,
$\sum|h(B_{n})|,\,\sum<1$
$A(1)\in(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}D(1)$
4.8 $v\neq 1,\infty$ $\xi\mathbb{R}_{\geq 0},D(1)=+\infty$ $A(1)=+\infty$
4.9 $v\neq 1,{}_{1}$, $\xi$-c. $\xi\mathbb{R}_{\geq 0},B(1)$
$\lim h(A_{n})=$
$(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}\in(0,1)$
4.10.1 $v\neq 1,{}_{1}$, $\xi$-c. $B(1)\neq 0,B_{v}(1),C_{v}(1)$ $\lim h(A_{n})\in(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}$
4.10.2 $v\neq 1,{}_{1}$, 0-c. $B(1)\neq 0,B_{v}(1),C_{v}(1)$ $A(1)\in(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}$
4.11.1 $v\neq 1,{}_{1}$, $\xi$-c.
$B(1),B_{v}(1),C_{v}(1)$,
$E(1)$
$\lim h(A_{n})\in(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}$
4.11.2 $v\neq 1,{}_{1}$, 0-c.
$B(1),B_{v}(1),C_{v}(1)$,
$E(1)$
$A(1)\in(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}$
4.12.1 $v\neq 1,{}_{2}$, $\xi$-c.
$B_{v}(1),\sum h(B_{n})<1$,
$\sum<1$
$\lim h(A_{n})\in(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}$
4.12.2 $v\neq 1,{}_{2}$, 0-c.
$B_{v}(1),\sum h(B_{n})<1$,
$\sum<1$
$A(1)\in(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}$
4.13.1 $v\neq 1,\infty$, $\xi$-c. $\xi\mathbb{R}_{\geq 0},B(1)=+\infty$ $\lim h(A_{n})=1$
4.13.2 $v\neq 1,\infty$, 0-c.
$\xi\mathbb{R},B(x),C_{v}(x)\in H_{1}$,
$B_{v}(1),B(1)=\pm\infty$
$\lim A(x)\in\{-1,1\}$
The first column lists the theorems. The second column lists the theorem types, with 1 and 2 abbreviating $\exists\,U(1)_{1}$ and $\exists\,U(1)_{2}$, respectively. By $\infty$ we abbreviate that $U(1)_{1}=\infty$. The third and fourth columns list the assumptions and conclusions of the theorems. We always omit indices h, 1/ξ, n→∞ and x↗1, and factor $\xi^{-n}$. We omit indices 1, 2, and v whenever they can be deduced from the context. By writing $A(1)$ in the third column and first row, we mean the assumption that $A_{h}(1)_{1}$ exists, and similarly in other cases. By writing $\xi\mathbb{R}_{\geq 0}$ and $\xi\mathbb{R}$ we mean the assumptions $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}$, respectively; else $h\colon\mathbb{N}_{2}\to\mathbb{C}$. $\sum|h(B_{n})|$ abbreviates $\sum_{n\geq 1}|h(B_{n})|$, and $\sum<1$ abbreviates the complicated bound assumed in the respective theorem.